21 Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make...

21

Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make and sell more product. The machine cost $1.34 million and create incremental cash flows of $670,775.00 each year for the next five years. The cost of capital is 11.07%. What is the net present value of the J-Mix 2000?

Answer format: Currency: Round to: 2 decimal places.

#2 Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make and sell more product. The machine cost $1.91 million and create incremental cash flows of $549,955.00 each year for the next five years. The cost of capital is 10.60%. What is the internal rate of return for the J-Mix 2000?

Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))

#3 Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make and sell more product. The machine cost $1.53 million and create incremental cash flows of $586,555.00 each year for the next five years. The cost of capital is 10.97%. What is the profitability index for the J-Mix 2000?

Answer format: Number: Round to: 3 decimal places.

#4 Derek plans to buy a $25,191.00 car. The dealership offers zero percent financing for 52.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 4.06%.

Answer format: Currency: Round to: 2 decimal places.

#5 Derek decides to buy a new car. The dealership offers him a choice of paying $600.00 per month for 5 years (with the first payment due next month) or paying some $28,855.00 today. He can borrow money from his bank to buy the car. What interest rate makes him indifferent between the two options?
Answer #1

1.
=-1340000+670775/11.07%*(1-1/1.1107^5)=1134756.07240093

2.
=RATE(5,549955,-1910000)=13.5205832901978%

3.
=1+(-1530000+586555/10.97%*(1-1/1.1097^5))/1530000=1.41796331197794

4.
=25191-PV(4.06%/12,52,-25191/52,0,1)=2049.16678130547

P.S.: I am not allowed to answer more than 4 questions